OpenStudy (nathanjhw):

Suppose you want to find an equation for a curve whose slope at any point (x, y) is x/(6y^2). You'd need to solve the differential equation dy/dx = x/(6y^2). Using separation of variables to solve this equation, you get (where C is an arbitrary constant of integration)

OpenStudy (nathanjhw):

6y^3= x^2 + C 3y^3 = 2x^2 + C 1/3y^3 = 1/2x^2 + C 2y^3 = x^2 + C 2y^3 = 1/2x^2 + C

OpenStudy (loser66):

What do you think?

OpenStudy (nathanjhw):

The last one @Loser66

OpenStudy (loser66):

yup